# Changing the Training Data Distribution to Reduce Simplicity Bias Improves In-distribution Generalization

### v2

@@ Abstract
Can we modify the training data distribution to encourage the underlying optimization method toward finding solutions with superior generalization performance on in-distribution data? In this work, we approach this question for the first time by comparing the inductive bias of gradient descent (GD) with that of sharpness-aware minimization (SAM). By studying a two-layer CNN, we rigorously prove that SAM learns different features more uniformly, particularly in early epochs. That is, SAM is less susceptible to simplicity bias compared to GD. We also show that examples containing features that are learned early are separable from the rest based on the model’s output. Based on this observation, we propose a method that (i) clusters examples based on the network output early in training, (ii) identifies a cluster of examples with similar network output, and (iii) upsamples the rest of examples only once to alleviate the simplicity bias. We show empirically that USEFUL effectively improves the generalization performance on the original data distribution when training with various gradient methods, including (S)GD and SAM. Notably, we demonstrate that our method can be combined with SAM variants and existing data augmentation strategies to achieve, to the best of our knowledge, state-of-the-art performance for training ResNet18 on CIFAR10, STL10, CINIC10, Tiny-ImageNet; ResNet34 on CIFAR100; and VGG19 and DenseNet121 on CIFAR10.
@@ 3.1 Theoretical Settings
Notation. We use lowercase letters, lowercase boldface letters, and uppercase boldface letters to denote scalars $(a)$, vectors $(\boldsymbol{v})$, and matrices $(\boldsymbol{W})$. For a vector $\boldsymbol{v}$, we use $\left\|\boldsymbol{v}\right\|_{2}$ to denote its Euclidean norm. Given two sequence $\{x_{n}\}$ and $\{y_{n}\}$, we denote $x_{n}=O(y_{n})$ if $|x_{n}|\leq C_{1}|y_{n}|$ for some absolute positive constant $C_{1}$, $x_{n}=\Omega(y_{n})$ if $|x_{n}|\geq C_{2}|y_{n}|$ for some absolute positive constant $C_{2}$, and $x_{n}=\Theta(y_{n})$ if $C_{3}|y_{n}|\leq|x_{n}|\leq C_{4}|y_{n}|$ for some absolute constant $C_{3},C_{4}>0$. Besides, we use $\tilde{O}(\cdot),\tilde{\Omega}(\cdot),$ and $\tilde{\Theta}(\cdot)$ to hide logarithmic factors in these notations. Furthermore, we denote $x_{n}=\text{poly}(y_{n})$ if $x_{n}=O(y_{n}^{D})$ for some positive constant D, and $x_{n}=\text{polylog}(y_{n})$ if $x_{n}=\text{poly}(\log(y_{n}))$.
@@
  • Random noise. The rest of $P-2$ patches are Gaussian noise $\boldsymbol{\xi}$ that are independently drawn from $N(0,(\sigma_{p}^{2}/d)\cdot\mathbf{I}_{d})$ with $\sigma_{p}$ as an absolute constant.
@@ 4 Method: UpSample Early For Uniform Learning (USEFUL)
[…]
@@ 5.1 Toy Datasets
Datasets. Following [18], our toy dataset consists of training and test sets, each containing 10K examples generated from the data distribution defined in 3.1 with dimension $d\!=\!50$ and $P\!=\!3$. We set $\beta_{e}\!=\!1,\beta_{d}\!=\!0.2,\alpha\!=\!0.9$, and $\sigma_{p}/\sqrt{d}\!=\!0.125$. We also consider a scenario with larger $\beta_{d}\!=\!0.4$. We shuffle the order of patches randomly to confirm that our theory holds with arbitrary order of patches.
@@ Appendix A Formal Proofs
Notation. In this paper, we use lowercase letters, lowercase boldface letters, and uppercase boldface letters to respectively denote scalars $(a)$, vectors $(\boldsymbol{v})$, and matrices $(\boldsymbol{W})$. For a vector $\boldsymbol{v}$, we use $\left\|\boldsymbol{v}\right\|_{2}$ to denote its Euclidean norm. Given two sequence $\{x_{n}\}$ and $\{y_{n}\}$, we denote $x_{n}=O(y_{n})$ if $|x_{n}|\leq C_{1}|y_{n}|$ for some absolute positive constant $C_{1}$, $x_{n}=\Omega(y_{n})$ if $|x_{n}|\geq C_{2}|y_{n}|$ for some absolute positive constant $C_{2}$, and $x_{n}=\Theta(y_{n})$ if $C_{3}|y_{n}|\leq|x_{n}|\leq C_{4}|y_{n}|$ for some absolute constant $C_{3},C_{4}>0$. In addition, we use $\tilde{O}(\cdot),\tilde{\Omega}(\cdot),$ and $\tilde{\Theta}(\cdot)$ to hide logarithmic factors in these notations. Furthermore, we denote $x_{n}=\text{poly}(y_{n})$ if $x_{n}=O(y_{n}^{D})$ for some positive constant D, and $x_{n}=\text{polylog}(y_{n})$ if $x_{n}=\text{poly}(\log(y_{n}))$.
@@
=-= - $\frac{3}{N}\sum_{i=1}^{N}l_{i}^{(t)}y_{i}\sum_{p=1}^{P}\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{x}^{(p)}\rangle^{2}\cdot\boldsymbol{x}^{(p)}$
=-= - $\frac{3}{N}\sum_{i=1}^{N}l_{i}^{(t)}\left(\beta_{d}^{3}\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{v}_{d}\rangle^{2}\boldsymbol{v}_{d}+\beta_{e}^{3}\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{v}_{e}\rangle^{2}\boldsymbol{v}_{e}+y_{i}\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{\xi}_{i}\rangle^{2}\boldsymbol{\xi}_{i}\right)-$
@@
Note that $10<l_{i}^{(t)}<$ 1 due to the property of the sigmoid function. Furthermore, we similarly consider that the sum of the sigmoid terms for all time steps is bounded up to a logarithmic dependence [11]. The sigmoid term is considered small for a $\kappa$ such that
@@
=-= - $\frac{3}{N}\sum_{i=1}^{N}l_{i,\boldsymbol{\epsilon}}^{(t)}y_{i}\sum_{p=1}^{P}\langle\boldsymbol{w}_{j,\boldsymbol{\epsilon}}^{(t)},\boldsymbol{x}^{(p)}\rangle^{2}\cdot\boldsymbol{x}^{(p)}$
=-= - $\frac{3}{N}\sum_{i=1}^{\alpha N}l_{i,\boldsymbol{\epsilon}}^{(t)}\left(\beta_{d}^{3}\langle\boldsymbol{w}_{j,\boldsymbol{\epsilon}}^{(t)},\boldsymbol{v}_{d}\rangle^{2}\boldsymbol{v}_{d}+\beta_{e}^{3}\langle\boldsymbol{w}_{j,\boldsymbol{\epsilon}}^{(t)},\boldsymbol{v}_{e}\rangle^{2}\boldsymbol{v}_{e}+y_{i}\langle\boldsymbol{w}_{j,\boldsymbol{\epsilon}}^{(t)},\boldsymbol{\xi}_{i}\rangle^{2}\boldsymbol{\xi}_{i}\right)-$
@@
00 0 $\leq\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{v}_{d}\rangle-\rho\leq\langle\boldsymbol{w}_{j,\boldsymbol{\epsilon}}^{(t)},\boldsymbol{v}_{d}\rangle\leq\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{v}_{d}\rangle$ (37)
@@ Appendix B Auxiliary Lemmas
Let $X\in\mathbb{R}^{d}$ be a Gaussian random vector, $X\sim\mathcal{N}(0,\sigma^{2}I_{d})$. Then with probability at least $1-o(1)$, we have $\|X\|_{2}^{2}=\Theta(\sigma^{2}\sqrt{d})$.
@@
Let $N=O(poly(d))$. The following hold with probability at least $1-o(1)$:
@@ D.7 Comparison with simplicity bias mitigation methods
[…]

### v3

# arXiv:2404.17768v3[cs.LG] 02 Mar 2026

@@ Abstract
Can we modify the training data distribution to encourage the underlying optimization method toward finding solutions with superior generalization performance on in-distribution data? In this work, we approach this question for the first time by comparing the inductive bias of gradient descent (GD) with that of sharpness-aware minimization (SAM). By studying a two-layer CNN, we rigorously prove that SAM learns different features more uniformly, particularly in early epochs. That is, SAM is less susceptible to simplicity bias compared to GD. We also show that examples containing features that are learned early are separable from the rest based on the model’s output. Based on this observation, we propose a method that (i) clusters examples based on the network output early in training, (ii) identifies a cluster of examples with similar network output, and (iii) upsamples the rest of examples only once to alleviate the simplicity bias. We show empirically that USEFUL effectively improves the generalization performance on the original data distribution when training with various gradient methods, including (S)GD and SAM. Notably, we demonstrate that our method can be combined with SAM variants and existing data augmentation strategies to achieve, to the best of our knowledge, state-of-the-art performance for training ResNet18 on CIFAR10, STL10, CINIC10, Tiny-ImageNet; ResNet34 on CIFAR100; and VGG19 and DenseNet121 on CIFAR10. Our code is available at https://github.com/BigML-CS-UCLA/TADA.
@@ 3.1 Theoretical Settings
[…]
@@
  • Random noise. The rest of P2P-2 patches are Gaussian noise $\boldsymbol{\xi}$ that are independently drawn from $N(0,(\sigma_{p}^{2}/d)\cdot\mathbf{I}_{d})$ with $\sigma_{p}$ as an absolute constant.
@@ 4 Method: UpSample Early For Uniform Learning (USEFUL)
Figure 3: GD (blue) vs. SAM (orange) on toy datasets. Data is generated based on Definition 3.1 with different $\beta_{d}$ and fixed $\beta_{e}=1,$ $\alpha=0.9$. $\cdot\cdot$ and -- lines denote the alignment (i.e., inner product) of fast-learnable ($\boldsymbol{v}_{e}$) and slow-learnable ($\boldsymbol{v}_{d}$) features with the model weight ($\boldsymbol{w}_{j}^{(t)}$). (a), (b) GD and SAM first learn the fast-learnable feature. Notably, GD learns the fast-learnable feature very early. (c) Test accuracy of GD & SAM improves by increasing the strength of the slow-learnable feature.
@@ 5.1 Toy Datasets
Datasets. Following [18], our toy dataset consists of training and test sets, each containing 10K examples generated from the data distribution defined in 3.1 with dimension d=50d\!=\!50 and P=3P\!=\!3. We set $\beta_{e}\!=\!1,\beta_{d}\!=\!0.2,\alpha\!=\!0.9$, and $\sigma_{p}/\sqrt{d}\!=\!0.125$. We also consider a scenario with larger $\beta_{d}\!=\!0.4$. We shuffle the order of patches randomly to confirm that our theory holds with arbitrary order of patches.
@@ Appendix A Formal Proofs
[…]
@@
$=$- $\frac{3}{N}\sum_{i=1}^{N}l_{i}^{(t)}y_{i}\sum_{p=1}^{P}\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{x}^{(p)}\rangle^{2}\cdot\boldsymbol{x}^{(p)}$
$=$- $\frac{3}{N}\sum_{i=1}^{N}l_{i}^{(t)}\left(\beta_{d}^{3}\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{v}_{d}\rangle^{2}\boldsymbol{v}_{d}+\beta_{e}^{3}\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{v}_{e}\rangle^{2}\boldsymbol{v}_{e}+y_{i}\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{\xi}_{i}\rangle^{2}\boldsymbol{\xi}_{i}\right)$-
@@
Note that $0<l_{i}^{(t)}<1$ due to the property of the sigmoid function. Furthermore, we similarly consider that the sum of the sigmoid terms for all time steps is bounded up to a logarithmic dependence [11]. The sigmoid term is considered small for a $\kappa$ such that
@@
$=$- $\frac{3}{N}\sum_{i=1}^{N}l_{i,\boldsymbol{\epsilon}}^{(t)}y_{i}\sum_{p=1}^{P}\langle\boldsymbol{w}_{j,\boldsymbol{\epsilon}}^{(t)},\boldsymbol{x}^{(p)}\rangle^{2}\cdot\boldsymbol{x}^{(p)}$
$=$- $\frac{3}{N}\sum_{i=1}^{\alpha N}l_{i,\boldsymbol{\epsilon}}^{(t)}\left(\beta_{d}^{3}\langle\boldsymbol{w}_{j,\boldsymbol{\epsilon}}^{(t)},\boldsymbol{v}_{d}\rangle^{2}\boldsymbol{v}_{d}+\beta_{e}^{3}\langle\boldsymbol{w}_{j,\boldsymbol{\epsilon}}^{(t)},\boldsymbol{v}_{e}\rangle^{2}\boldsymbol{v}_{e}+y_{i}\langle\boldsymbol{w}_{j,\boldsymbol{\epsilon}}^{(t)},\boldsymbol{\xi}_{i}\rangle^{2}\boldsymbol{\xi}_{i}\right)$-
@@
$0$ $\leq\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{v}_{d}\rangle-\rho\leq\langle\boldsymbol{w}_{j,\boldsymbol{\epsilon}}^{(t)},\boldsymbol{v}_{d}\rangle\leq\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{v}_{d}\rangle$ (37)
@@ Appendix B Auxiliary Lemmas
Let $X\in\mathbb{R}^{d}$ be a Gaussian random vector, $X\sim\mathcal{N}(0,\sigma^{2}I_{d})$. Then with probability at least 1o(1)1-o(1), we have $\|X\|_{2}^{2}=\Theta(\sigma^{2}\sqrt{d})$.
@@
Let $N=O(poly(d))$. The following hold with probability at least 1o(1)1-o(1):
@@ D.7 Comparison with simplicity bias mitigation methods
While previous works show that reducing simplicity bias benefits the OOD settings, we show that reducing the simplicity bias also benefits the ID settings. To confirm our hypothesis on our simplicity bias mitigation baselines, we applied EIIL [14] and JTT [48] to train ResNet18 on CIFAR10. The choice of the baselines is because they have publicly available code and fewer hyperparameters to tune in our limited rebuttal time. For EIIL, we tuned lr $\in${1e-1, 1e-2, 1e-3, 5e-4, 1e-4}, number of epochs $\in${1, 2, 4, 8} for training the reference model, and the weight decay $\in${1e-3, 5e-4, 1e-4} for training GroupDRO. For JTT, we tuned the separating epoch $\in${4, 5, 6, 7, 8, 10} and upsampling factor $\in${2, 3, 4, 5}, while lr and weight decay follow the standard training of ResNet18 on CIFAR10. Table 9 shows that all methods successfully reduce the simplicity bias, yielding an improvement over SGD. While EIIL requires tuning 3 hyperparameters with a total of 60 combinations, USEFUL only requires one hyperparameter, which is the separating epoch within a small range (around the time when the slope of the training loss curve diminishes).